OpenStudy (anonymous):

find the equation of an altitude from A in triangle ABC. A(2,8) B(-6,-4) C(8,0) thank you!

jimthompson5910 (jim_thompson5910):

First find the slope of the line through points B and C

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

-8,-12 and 6,-8

jimthompson5910 (jim_thompson5910):

how are you getting that

jimthompson5910 (jim_thompson5910):

the slope is just a single number btw

OpenStudy (anonymous):

idk...

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

so in this case, you would get m = (y2 - y1)/(x2 - x1) m=(0-(-4))/(8-(-6)) m = (0+4)/(8+6) m = 4/14 m = 2/7

OpenStudy (anonymous):

2/3 and -.75

jimthompson5910 (jim_thompson5910):

That's the slope of the line through B and C

jimthompson5910 (jim_thompson5910):

see how I got that?

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

so what's the perpendicular slope

OpenStudy (anonymous):

ya but what about the altitude?

jimthompson5910 (jim_thompson5910):

we're getting there (slowly but surely) what's the perpendicular slope

OpenStudy (anonymous):

7/2?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

but you have to change the sign as well, so it's negative

jimthompson5910 (jim_thompson5910):

so we know this the altitude has a slope of -7/2 the altitude goes through the point (2,8) do you see how to use this info to find the equation?